OpenStudy (anonymous):

A direct variation function contains the points (2, 14) and (4, 28). Which equation represents the function?

OpenStudy (anonymous):

Plug them into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{28-14}{4-2}\) Subtract, what's 28 - 14?

OpenStudy (anonymous):

14